# Agent Control Protocol ACP v1.19: Admission Control for Agent Actions Technical Specification and Reference Implementation

### 25 Mar 2026 v5

@@ 12.1  Evaluation Goals
  • Q4: Does ACP maintain admission control effectiveness under adversarial patterns, including cooldown evasion and distributed multi-agent attacks?
@@ 12.2  Experimental Setup
All experiments use the Go reference implementation (v1.19) on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, GOMAXPROCS=8). The evaluation uses three components: the stateless evaluation function (pkg/risk.Evaluate), an in-memory state backend (InMemoryQuerier), and the ACR-1.0 compliance runner (Section 13). Production deployments with persistent storage backends will incur additional I/O latency not measured here.
@@ 12.3  Q1: Computational Overhead
Table 10 reports go test -bench results. The full evaluation path (all three F_anom rules active) completes in 921 ns. The cooldown short-circuit path (88 ns) bypasses the risk function entirely.
Answer to Q1: The admission check adds between 88 ns and 921 ns of latency per request. Given that the actions ACP governs (API calls, financial transactions, file mutations) operate in the millisecond-to-second range, the overhead is at least three orders of magnitude below the cost of the governed action.
@@ 12.4  Throughput and State Backend Contention
Table 11: ACP throughput under increasing concurrency (Intel Core i7-8665U, Go 1.22, Windows/amd64).
ACP sustains approximately 920,000 requests per second at 10 concurrent workers. Throughput degrades gradually as concurrency increases, reaching approximately 712,000 requests per second at 500 workers. Degradation is gradual rather than catastrophic, and no correctness violations were observed under load.
@@ 12.6  Q3: Reproducibility and External Verifiability
12.7  Adversarial Evaluation (ACP-RISK-2.0)
We evaluate ACP under the adversary model defined in Section 16. The following attack strategies instantiate specific adversary capabilities, including adaptive request patterns and multi-agent coordination. We conducted three targeted experiments using the same LedgerQuerier abstraction and execution contract as the ACR-1.0 compliance runner, addressing adversarial robustness and real-world state-backend scaling.
@@ Results.
ACP triggered cooldown after exactly 3 real DENIED decisions (requests #0, #2, #4). All subsequent requests—regardless of risk score—were short-circuited to COOLDOWN_ACTIVE at Step 2 of the evaluation pipeline ($\sim$88 ns, see Table 10).
@@
Interpretation. Per-agent admission control provides strong containment for single-agent threats. For distributed attacks, an attacker with $N$ distinct agent identities can execute $3N$ high-risk requests before all agents are individually blocked. Mitigating this requires cross-agent attribution at the policy layer (e.g., shared-resource rate limiting, identity clustering), which is deliberately outside the scope of ACP’s per-agent admission model. This is a design boundary, not a correctness failure.
@@
Setup. 500 agents execute 20 requests each (10,000 total, fully concurrent) against two backend implementations: the reference InMemoryQuerier (Go mutex over in-process slices) and a minimal RedisQuerier (Redis 7 sorted sets via go-redis/v9, Docker loopback, one Redis command per ledger operation, no pipelining). This scenario exercises the adversary’s ability to generate high request volumes and exploit shared state contention.
@@
Backend Duration Throughput
InMemoryQuerier $\sim$28 ms $\sim$350,000 req/s
RedisQuerier $\sim$4.6 s $\sim$2,100 req/s
Table 16: Experiment 3: State Backend Stress (500 agents $\times$ 20 requests =10,000=10{,}000 total). InMemoryQuerier variance is high under concurrent load ($\pm$30%) due to Go mutex scheduling; RedisQuerier is stable ($\pm$4%).
Interpretation. The lower throughput of InMemoryQuerier under adversarial load ($\sim$350k req/s) compared to the 500-worker baseline of $\sim$712k req/s (Table 11) is expected: the adversarial workload generates significantly higher contention because every request updates multiple sorted lists under a shared mutex. RedisQuerier eliminates the global Go mutex entirely, trading it for per-operation network RTT. A production-grade Redis backend using command pipelining or server-side Lua scripts would reduce latency to a single round-trip per request. The key architectural result remains: ACP protocol semantics are stable across backends; deployment performance scales with the quality of the state backend, not with the admission control logic itself.
@@
Attack 1 (cooldown evasion via pattern alternation) directly targets the cooldown mechanism. Results show that cooldown activation cannot be indefinitely avoided: adversarial alternation delays but does not prevent enforcement. This confirms that cooldown is eventually enforced under repeated violations.
@@
Attack 2 (distributed multi-agent attack) targets system-wide coordination. Results show that ACP enforces constraints independently per agent, maintaining isolation. However, coordinated behavior across agents is not detected, which constitutes an explicit design limitation rather than a violation.
@@
Summary. The evaluated attacks confirm that ACP enforces its core security properties under adversarial conditions. No violations are observed; instead, limitations arise from explicit design trade-offs, particularly in cross-agent correlation. These results reinforce the central design insight: ACP is compute-cheap but state-sensitive.

### 30 Apr 2026 v10

@@ 13.1  Evaluation Goals
  • Q4: Does ACP maintain admission control effectiveness under adversarial patterns, including cooldown evasion, distributed multi-agent attacks, and token replay?
@@ 13.2  Experimental Setup
All experiments use the Go reference implementation (v1.30) on an Intel Core i7-8665U @ 1.90GHz (Go 1.22 [19], GOMAXPROCS=8). The evaluation uses three components: the stateless evaluation function (pkg/risk.Evaluate), an in-memory state backend (InMemoryQuerier), and the ACR-1.0 compliance runner (Section 14). Production deployments with persistent storage backends will incur additional I/O latency not measured here.
@@ 13.3  Q1: Computational Overhead
Table 12 reports go test -bench results. The full evaluation path (all three F_anom rules active) completes in 832 ns. The cooldown short-circuit path (78 ns) bypasses the risk function entirely.
Answer to Q1: The admission check adds between 78 ns and 832 ns of decision latency per request (p50). For single-goroutine CPU-bound microbenchmarks (InMemoryQuerier, no I/O), variance is negligible: p99 is within 5–8% of the reported p50 values (p99 $\lesssim$ 900 ns for full evaluation; p99 $\lesssim$ 85 ns for the cooldown path). In production deployments, latency variance is dominated by state backend I/O (Redis RTT: 0.1–1 ms; PostgreSQL: 1–5 ms), not by evaluation logic. Given that the actions ACP governs (API calls, financial transactions, file mutations) operate in the millisecond-to-second range, the overhead is at least three orders of magnitude below the cost of the governed action. Note that this measures decision-logic latency only; aggregate throughput is governed by state backend access patterns and is reported separately in the following subsection.
@@ 13.4  Throughput and State Backend Contention
Table 13: ACP throughput under increasing concurrency (Intel Core i7-8665U, Go 1.22, Windows/amd64, InMemoryQuerier). This is a concurrent pipeline benchmark distinct from the go test -bench microbenchmarks in Table 12; the 1,720,000 req/s figure in Table 12 reflects the engine-only evaluation path with a single goroutine.
ACP sustains approximately 920,000 requests per second at 10 concurrent workers (pipeline benchmark; Table 13). Throughput degrades gradually as concurrency increases, reaching approximately 712,000 requests per second at 500 workers. Degradation is gradual rather than catastrophic, and no correctness violations were observed under load.
@@ 13.6  Q3: Reproducibility and External Verifiability
13.7  Adversarial Evaluation (ACP-RISK-2.0 and ACP-RISK-3.0)
We evaluate ACP under the adversary model defined in Section 19. The following attack strategies instantiate specific adversary capabilities, including adaptive request patterns and multi-agent coordination. We conducted five targeted experiments using the same LedgerQuerier abstraction and execution contract as the ACR-1.0 compliance runner, addressing adversarial robustness, real-world state-backend scaling, and cryptographic-layer replay resistance.
@@ Results.
ACP triggered cooldown after exactly 3 real DENIED decisions (requests #0, #2, #4). All subsequent requests—regardless of risk score—were short-circuited to COOLDOWN_ACTIVE at Step 2 of the evaluation pipeline ($\sim$78 ns, see Table 12).
@@
Interpretation. Per-agent admission control provides strong containment for single-agent threats. For distributed attacks, an attacker with $N$ distinct agent identities can execute $3N$ high-risk requests before all agents are individually blocked. Mitigating this requires cross-agent attribution at the policy layer (e.g., shared-resource rate limiting, identity clustering), which is deliberately outside the scope of ACP’s per-agent admission model. This behavior reflects a structural limitation of per-agent enforcement rather than a correctness failure of the protocol. The implications of this design are discussed in §13.8 (Distributed Attack Surface).
@@
Setup. 500 agents execute 20 requests each (10,000 total, fully concurrent) against three backend implementations: (1) InMemoryQuerier (Go mutex over in-process slices); (2) RedisQuerier (Redis 7 [38] sorted sets via go-redis/v9, Docker loopback, one Redis command per ledger operation, $\sim$7–8 RTTs per request); (3) RedisPipelinedQuerier (two Redis pipeline round-trips per request: one for all 4 reads before Evaluate, one for all writes after). RedisPipelinedQuerier is implemented in compliance/adversarial/redis_pipelined.go. This scenario exercises the adversary’s ability to generate high request volumes and exploit shared state contention.
@@
Backend RTTs/req Throughput (req/s) Duration Relative
[…]
Key insight. These results reinforce a central property of ACP: the protocol is compute-cheap but state-sensitive. Admission decisions incur minimal computational cost, while system performance is primarily determined by the efficiency of state access and synchronization. ACP does not eliminate the cost of distributed state access—it makes that cost explicit and measurable.
Experiment 4: Controlled Latency Injection
Hypothesis. System throughput is bounded by state backend latency, not by the admission control logic itself. This experiment isolates that effect by injecting fixed per-call latency into the LedgerQuerier interface without involving external infrastructure.
Setup. To characterize admission throughput as a function of backend latency independently of specific infrastructure, we evaluate ACP using a DelayedQuerier wrapper that injects a fixed per-call latency into the LedgerQuerier interface. Each Evaluate() call in the non-cooldown path triggers four LedgerQuerier calls (CooldownActive, CountRequests, CountDenials, CountPattern); total per-request injected latency is approximately $4\times$ the per-call value. This experiment injects controlled latency and does not model full network variability or distributed system effects.
Injected latency/call Total latency/req Throughput (req/s) Scenario
0 µs (InMemoryQuerier) $\sim$0 µs $\sim$920,000 Process-local baseline
250 µs $\sim$1 ms $\sim$4,100 Fast datacenter
1 ms $\sim$4 ms $\sim$1,600 Busy backend
5 ms $\sim$20 ms $\sim$470 Cross-datacenter
Table 19: Controlled latency injection via DelayedQuerier (10 workers, Intel i7-8665U, Go 1.22). Throughput degrades proportionally with injected backend latency; the Evaluate() decision function contributes 739–832 ns per request in all configurations. The LedgerQuerier abstraction enables this separation.
Throughput decreases proportionally with injected latency, supporting the claim that the admission control logic is not the performance bottleneck. The Evaluate() function contributes a fixed 739–832 ns per request (p50; Table 12) regardless of backend latency; system throughput is dominated by state access patterns. This validates the LedgerQuerier abstraction as the correct boundary for deploying ACP with heterogeneous backends (in-process, Redis, CockroachDB, or geo-replicated stores) without modifying protocol semantics.
Experiment 5: Token Replay Attack
[…]
Results.
Case Requests ESCALATED DENIED Cooldown-blocked
Normal baseline (Case 1) 10 0 0 0
Sequential replay (Case 2) 10 3 3 4
Concurrent replay (Case 3) 20 3 3 14
Near-identical replay (Case 4) 10 10 0 0
[…]
For sequential replay, the RS trajectory is: requests 1–3 at RS =55=55 (ESCALATED), requests 4–6 at RS =70=70 (DENIED; Rule 3 triggered), requests 7–10 as COOLDOWN_ACTIVE. Figure 1 plots the RS per request for normal vs. replay traffic. Concurrent replay (Case 3) produces the same trajectory: InMemoryQuerier’s mutex serializes LedgerQuerier reads, so concurrency does not bypass accumulation; 14/20 requests are blocked.
Interpretation. ACP converts identical replay attempts into observable state transitions via pattern accumulation. F_anom Rule 3 amplifies the RS of a borderline-ESCALATED token to DENIED after $Y=3$ occurrences within the 5-minute window, without any nonce-tracking logic. Near-identical variation (Case 4) evades Rule 3 by producing distinct patternKeys, demonstrating that replay resistance is bounded: tokens with varied payload fields and moderate base RS stay below the DENIED threshold. This constitutes an explicit design boundary rather than a correctness failure; see §13.8.
1122334455667788991010020405570100DENIED (>>69)APPROVED ($\leq$39)COOLDOWNRule 3Request indexRisk score (RS)Normal trafficReplay attack
Figure 1: Replay vs. normal traffic: RS per request (Case 2 sequential replay). Requests 1–3: RS =55=55 (ESCALATED). Request 4: F_anom Rule 3 fires (+15 RS) $\to$ RS =70=70 (DENIED). Request 7+: COOLDOWN_ACTIVE (sentinel 105). Normal traffic stays at RS =0=0 throughout.
@@
Experiment 6: Stateless vs. Stateful Admission
Hypothesis. Stateless policy evaluation cannot detect temporal abuse patterns that require accumulation over execution traces. A stateful admission model (ACP), incorporating temporal signals such as anomaly accumulation and cooldown, can detect and block such patterns under identical scoring conditions.
Setup. We compare ACP against a stateless variant that reuses the exact same scoring function and policy thresholds but omits all stateful signals. The stateless engine is implemented by invoking the same Evaluate() function with a NullQuerier, which returns zero for all state queries (request counts, denials, pattern counts, and cooldown state). This ensures that both engines share identical logic, differing only in the presence of state.
The workload consists of 500 sequential requests from a single agent with:
  • Capability: acp:cap:financial.transfer
  • Resource class: public
Under the ACP scoring model, this yields a base risk score $RS_{base}=35$, which is below the approval threshold ($ApprovedMax=39$). Therefore, each request is individually valid and always approved by the stateless engine.
To isolate temporal accumulation effects, all requests are evaluated using a fixed timestamp. This ensures that all observations fall within the anomaly detection windows and removes time-based decay as a confounding variable. State updates (request count and pattern tracking) are recorded prior to each evaluation, capturing every attempted action regardless of outcome.
Results. Table 21 summarizes the primary outcome.
Table 21: Experiment 6 — Stateless vs. Stateful Admission
Engine Auto. Exec. Non-Approved Early Signal Enforcement Latency
Stateless 100% (500/500) 0% $\infty$ $\infty$
ACP (stateful) 0.4% (2/500) 99.6% (498/500) 3 actions 11 actions
Automatic execution measures the fraction of requests executed without human intervention. Non-approved includes ESCALATED, DENIED, and COOLDOWN outcomes.
Table 22 provides a detailed breakdown of ACP decisions.
Table 22: Experiment 6 — Decision Breakdown (500 requests)
Engine APPROVED ESCALATED DENIED COOLDOWN
Stateless 500 0 0 0
ACP 2 8 3 487
Detection Dynamics. ACP exhibits a two-stage response:
  • Early signal: the first ESCALATED decision occurs at request #3, when pattern-based anomaly detection (Rule 3) activates.
  • Enforcement: the first DENIED decision occurs at request #11, when combined anomaly signals (Rules 1 and 3) raise the risk score to the denial threshold.
  • Containment: after three denials, a cooldown is triggered at request #13, blocking all subsequent requests.
Figure 2 illustrates the temporal evolution of decisions under both models. The figure highlights that ACP enforcement emerges from accumulation over time, rather than from any individual request exceeding a threshold.
Insight. The stateless engine approves all 500 requests because each request is individually valid. It cannot distinguish between isolated actions and coordinated behavioral patterns. This result generalizes beyond the synthetic baseline: any stateless policy engine—including production systems such as Open Policy Agent (OPA) [36] and Cedar [2]—would produce the same outcome by construction, since they evaluate each request against a policy snapshot without access to the agent’s execution history. Per-request correctness is a necessary but not sufficient condition for behavioral safety under sustained agent operation.
In contrast, ACP detects the pattern after 3 actions and enforces blocking after 11 actions, limiting autonomous execution to 2 out of 500 requests.
The difference is not in scoring, but in the domain of evaluation: stateless systems evaluate requests, while ACP enforces constraints over execution traces (Definition 1).
Implication. This result demonstrates that temporal admission control cannot be reduced to per-request policy evaluation without loss of security properties. Systems that lack stateful enforcement admit adversarial sequences that are indistinguishable from benign behavior when evaluated in isolation.
5050100100150150200200250250300300350350400400450450500500APPROVEDESCALATEDDENIEDCOOLDOWN#3#11#13Request indexDecision stateStatelessACP (stateful)
Figure 2: Decision evolution under 500 repeated valid requests. The stateless engine approves all. ACP escalates after 3 actions, denies after 11 actions, and enters cooldown after 13 actions, blocking all subsequent requests.
Experiment 7: State-Mixing Vulnerability
Hypothesis. Rule 1 (high request rate) counts requests keyed by agentID only, without regard to capability or resource context. As a result, high-volume low-risk activity in one capability context can prime the Rule 1 counter and cause incorrect denial decisions in a subsequent, unrelated capability context. Rule 3 (repeated pattern), keyed by PatternKey(agentID, capability, resource), does not share this property.
Setup. We evaluate a single agent (legitimate-agent-1) across two sequential capability contexts using a shared InMemoryQuerier:
  • Control: fresh state; one financial.transfer/sensitive request. Expected RS=35+15=50=35+15=50 (ESCALATED).
  • Phase 1: 11 sequential data.read/public requests ($RS_{base}=0$). Each is individually approved. The requests increment the agentID-scoped Rule 1 counter, priming it past the threshold ($N=10$) without triggering any denial.
  • Phase 2: one financial.transfer/sensitive request with the shared state from Phase 1. Rule 1 fires (+20+20) because CountRequests(agentID, 60s) includes all prior requests regardless of context. Rule 3 does not fire because the pattern key is distinct from Phase 1.
Results.
Table 23: Experiment 7 — State-Mixing: RS and Decision Under Shared State
Condition Prior requests RS F_anom Rule 1 Decision
Control (clean state) 0 50 0 × ESCALATED
Phase 2 (contaminated) 11 70 20 DENIED
The RS elevation of +20+20 is attributable entirely to Rule 1 cross-context contamination. Rule 2 does not trigger (no prior denials in the shared state). Rule 3 does not trigger (the data.read/public pattern key is disjoint from financial.transfer/sensitive), confirming that the vulnerability is scoped to the agentID-level counter, not to the pattern-based mechanism.
Contamination threshold. The minimum number of requests in any context required to prime Rule 1 is $N+1=11$ (with $N=10$ per DefaultPolicyConfig). Below this threshold, cross-context contamination does not occur; above it, any legitimate high-volume workload can incorrectly escalate a subsequent admission decision.
Implication. This result demonstrates that ACP-RISK-2.0, as specified, does not isolate anomaly state per capability context. An agent that legitimately generates more than $N$ requests in one capability context within 60 seconds will receive an inflated risk score in all subsequent capability contexts within the same window. The effect is not detectable from the specification alone: it requires instantiating the engine across two distinct capability contexts and observing that CountRequests and CountPattern have different keying strategies.
This constitutes an explicit design boundary rather than an implementation defect; context isolation of Rule 1 requires per-capability request tracking (e.g., keying CountRequests by (agentID, capability)) at the cost of increased storage and reduced cross-context abuse detection. The trade-off and the implemented fix are discussed in §13.8 (State-Mixing Vulnerability).
Experiment 8: Context-Scoped Anomaly Fix (ACP-RISK-3.0)
Hypothesis. Scoping rate-based anomaly signals to interaction context eliminates cross-context interference while preserving detection of repeated behavior within each context.
Setup. We reproduce the state-mixing scenario from Experiment 7 under identical conditions, comparing ACP-RISK-2.0 and ACP-RISK-3.0. The workload consists of two phases: (1) a priming phase with 11 repeated low-risk requests in one context, and (2) a high-value request in a distinct context.
Results.
Table 24: Experiment 8 — Context-Scoped Anomaly Enforcement
Scenario Model RS Decision
Clean context RISK-2.0 50 ESCALATED
Clean context RISK-3.0 50 ESCALATED
State-mixing attack RISK-2.0 70 DENIED
State-mixing attack RISK-3.0 50 ESCALATED
Repeated single-context (11 req) RISK-2.0 85 DENIED
Repeated single-context (11 req) RISK-3.0 85 DENIED
Analysis. ACP-RISK-2.0 exhibits cross-context interference: activity in one context elevates risk in unrelated contexts via agent-level aggregation, producing a false denial in a legitimate context (RS $50\to 70$). ACP-RISK-3.0 eliminates this effect by scoping rate signals to the interaction context.
Repeated behavior within a single context continues to trigger denial after the same threshold (11 requests, RS =85=85), demonstrating that enforcement strength is preserved.
Conclusion. Context-scoped anomaly aggregation removes unintended cross-context coupling without weakening enforcement against repeated or high-frequency behavior.
@@
Attack 1 (cooldown evasion via pattern alternation) directly targets the cooldown mechanism. Results show that cooldown activation cannot be indefinitely avoided: adversarial alternation delays but does not prevent enforcement. This demonstrates that cooldown is eventually enforced under repeated violations.
@@
Attack 2 (distributed multi-agent attack) targets system-wide coordination. Results show that ACP applies constraints independently per agent, maintaining isolation under the experimental conditions. Coordinated behavior across agents is not detected, which constitutes an explicit design limitation rather than a violation.
@@
Bounded Replay Resistance.
Attack 4 (token replay) targets the cryptographic layer, specifically the absence of nonce tracking in ACP’s stateless admission model. Results show that identical replays are detected via F_anom Rule 3 pattern accumulation, which escalates borderline-ESCALATED tokens to DENIED after $Y=3$ occurrences in 5 minutes. Near-identical tokens with varied resource fields evade Rule 3, establishing an explicit boundary: ACP converts identical replays into observable state transitions but does not provide strong cryptographic replay prevention. This is empirically validated; see Section 13.8 for the corresponding Limitations note. The property is summarized: replay resistance is enforced through temporal validity and anomaly accumulation, not nonce-based prevention.
Summary. The evaluated attacks provide empirical evidence that ACP resists the tested adversarial scenarios. No violations are observed; instead, limitations arise from explicit design trade-offs, particularly in cross-agent correlation. These results reinforce the central design insight: ACP is compute-cheap but state-sensitive.
@@ 13.8  Limitations
Replay resistance is bounded. ACP does not implement nonce-based replay prevention; replay resistance is enforced through temporal validity (TimestampDrift flag) and anomaly accumulation (F_anom Rule 3). Tokens with varied resource fields and moderate base RS << DENIED threshold can evade pattern-based detection while generating ESCALATED outcomes, preventing cooldown activation (Experiment 5, Case 4). Nonce tracking requires cryptographic session state that is outside the scope of ACP’s stateless per-request admission model. This is an explicit design boundary, empirically characterized in Experiment 5. ACP provides graduated replay resistance across conformance levels: L1 enforces temporal validity (TimestampDrift); L2 adds anomaly accumulation (F_anom); L3 adds complete ledger auditing (ACP-LEDGER-1.3). Strict cryptographic nonce-based prevention is identified as a future specification extension, motivated by the ACP-LEDGER-1.0 hash-chained log work item.
State-Mixing Vulnerability in ACP-RISK-2.0 (fixed in v1.22 via ACP-RISK-3.0). ACP-RISK-2.0 aggregates anomaly signals at the agent level: F_anom Rules 1 and 2 are computed via CountRequests(agentID) and CountDenials(agentID), both indexed exclusively by agentID. This introduces cross-context interference: semantically distinct interaction patterns share a common anomaly accumulation register. This vulnerability is empirically demonstrated in Experiment 7, where 11 sequential data.read/public requests (each individually approved) prime the Rule 1 counter, causing a subsequent financial.transfer/sensitive request to receive an anomaly penalty of +20+20 and transition from ESCALATED (RS =50=50) to DENIED (RS =70=70) — a decision outcome that does not occur in clean state.
Attack construction. Consider agent $A_{1}$ sending three requests under context $\langle\textsf{read},\,\textsf{public}\rangle$ (BaseRisk = 0) followed by one request under $\langle\textsf{write},\,\textsf{sensitive}\rangle$ (BaseRisk = 50), with FLOOD_THRESHOLD = 3. Table 25 shows the evaluation under the current model, where pattern_count is indexed by agentID only.
Step Context pattern_count[A1] BaseRisk $F_{\textsf{anom}}$ Decision
1 (read, public) 1 0 0 approved
2 (read, public) 2 0 0 approved
3 (read, public) 3 (= threshold) 0 0 approved
4 (write, sensitive) 3 (inherited) 50 +25 denied
Table 25: State-mixing attack trace (baseline model). At Step 4, pattern_count[A1] = 3 is inherited from the (read, public) context. The anomaly penalty of +25 fires incorrectly: the (write, sensitive) request is a new context with zero prior occurrences, yet it is evaluated as if at the flood threshold. Total RS = 50 + 25 = 75 $\to$ denied.
Under a context-isolated model, pattern_count[A1][write/sensitive] = 0 at Step 4: $F_{\textsf{anom}}$ = 0, RS = 50, and the request is evaluated solely on its own context history. The cross-context penalty is eliminated.
Asymmetric keying. Notably, F_anom Rule 3 (repeated-pattern detection) already applies a per-context key: $\texttt{PatternKey}=\texttt{SHA-256}(\textit{agentID}\parallel\textit{capability}\parallel\textit{resource})$. Rules 1 and 2, however, retain agent-level granularity. This asymmetry means that pattern frequency is isolated per context while aggregate frequency and denial rate are not. This is not a limitation of pattern detection (Rule 3, which is correctly partitioned), but a consequence of inconsistent state partitioning across anomaly signals, leading to violations of context-consistent temporal enforcement.
Formal property. A corrected model enforces ContextIsolation:
For all agents $a$ and distinct contexts $\langle c_{1},r_{1}\rangle\neq\langle c_{2},r_{2}\rangle$, the anomaly state accumulated under $\langle c_{1},r_{1}\rangle$ does not contribute to the risk score computed for $\langle c_{2},r_{2}\rangle$.
The dual violation property StateMixingViolation asserts that a trace exists in which the decision for context $\langle c_{2},r_{2}\rangle$ is influenced by anomaly accumulation from $\langle c_{1},r_{1}\rangle$. In the current model (Rules 1 and 2 keyed by agentID only), StateMixingViolation holds: the trace in Table 25 is a concrete counterexample. In the extended model (all rules keyed by context), ContextIsolation is expected to hold by construction: each counter is indexed by $\texttt{SHA-256}(\textit{agentID}\parallel\textit{capability}\parallel\textit{resource})$, eliminating any shared mutable state between contexts. The D4 liveness property (EventuallyRejected) is structurally preserved: an agent flooding a single context still accumulates anomaly signals within that context until enforcement triggers. Formal TLC verification of the extended model is identified as future work. This connects to Schneider’s security automaton model [42]: a correct reference monitor must enforce temporal properties over equivalence classes of interactions; conflating distinct contexts violates that requirement.
Impact. The attack does not bypass enforcement entirely, but degrades anomaly accumulation, delaying or preventing escalation under mixed-context interaction patterns. An agent flooding a single context is still eventually enforced within that context; the vulnerability is that cross-context noise corrupts the signal for unrelated contexts, producing false positives (unwarranted penalties) or false negatives (missed escalation when benign context activity dilutes aggregate counters).
Cooldown leakage. Cooldown enforcement is also keyed at the agent level (CooldownActive(agentID)), allowing cross-context propagation of throttling effects. An agent whose cap-A activity triggers cooldown is blocked across all capabilities, including cap-B operations that have no prior violations. This constitutes a second cross-context interference vector, in the over-enforcement direction.
Mitigation path. The structural fix is a refinement of the state model: extend the per-context key used by Rule 3 to Rules 1 and 2, so all anomaly counters are indexed by $\texttt{SHA-256}(\textit{agentID}\parallel\textit{capability}\parallel\textit{resource})$. Cooldown granularity should similarly be evaluated per context for independent capabilities. This aligns the anomaly accumulation granularity with the semantic unit of the request, at the cost of requiring finer-grained ledger storage. This fix is implemented in ACP-RISK-3.0 and empirically validated in Experiment 8 (Section 13.7). Deployments relying on heterogeneous capability usage within a single agent identity should upgrade to ACP-RISK-3.0.
ACP-RISK-3.0 definition. ACP-RISK-3.0 refines anomaly aggregation by scoping rate-based signals to interaction context while preserving global denial history. Rule 1 is redefined to use context-scoped pattern counts:
$F_{\text{anom}}=\begin{cases}+20&\text{if }\mathrm{CountPattern}(k,\,60\,\text{s})>N\\ +15&\text{if }\mathrm{CountDenials}(a,\,24\,\text{h})\geq X\\ +15&\text{if }\mathrm{CountPattern}(k,\,5\,\text{min})\geq Y\end{cases}$ (1)
where $a$ is the agent identifier and $k=H(a\parallel\mathit{cap}\parallel\mathit{res})$ is the context key derived from agent, capability, and resource. Rule 2 (denial history) remains agent-scoped: cross-context denial history is a valid enforcement signal. Rule 3 is unchanged; it was already context-scoped. No changes to the LedgerQuerier interface are required.
Bounded Distributed Evasion.
ACP-RISK-3.0 does not enforce global rate limits across heterogeneous contexts. An adversary distributing requests across multiple contexts below the contextual rate threshold will trigger escalation via Rule 3 (after three repetitions per context) but not immediate denial. This behavior reflects the design objective of preventing autonomous execution rather than enforcing universal blocking: any repeated behavior in any context requires human oversight before proceeding.
Distributed Attack Surface. ACP-RISK-2.0 enforces anomaly detection and cooldown policies at the granularity of individual agent identities. As a result, anomaly signals such as request rate (Rule 1) and denial count (Rule 2) are accumulated independently per agent.
This design introduces a limitation under multi-agent adversarial scenarios. An attacker controlling $N$ distinct agent identities can distribute requests across agents, avoiding per-agent thresholds while still executing high-risk actions.
Quantified impact. Under the default policy (cooldown triggered after 3 denials per agent), an attacker can execute up to $3N$ high-risk requests before all agents are individually blocked. For example, with N=1, [^0] N=1{,}000 agents, up to 3,000 high-risk actions may be executed before full enforcement is reached. This behavior is empirically characterized in Experiment 2 (Table 17).
Root cause. The limitation arises from the absence of cross-agent aggregation in anomaly signals. Current rules are keyed by agentID, and no shared state is maintained across agents interacting with the same resource or system boundary.
Proposed extension. A mitigation is to introduce resource-level aggregation of anomaly signals. This requires extending the LedgerQuerier interface to support shared counters keyed by resource (or resource class). Conceptually, the aggregated denial signal would be expressed as:
$\texttt{globalDenials}(\mathit{resourceID},\,w)=\sum_{\mathit{agentID}}\texttt{denials}(\mathit{agentID},\,w)$
where $w$ is a sliding time window. When the aggregated signal exceeds a threshold, enforcement (e.g., resource-level cooldown) is applied across all agents.
Trade-offs. This approach requires a shared state backend and may introduce false positives in high-concurrency legitimate workloads. It also reduces isolation between independent agents and shifts part of the enforcement logic to the policy or orchestration layer.
Scope. ACP v1.30 enforces guarantees at the per-agent level. Cross-agent coordination is intentionally outside the scope of the core protocol and is expected to be implemented at the policy or orchestration layer when required by the deployment context. A future extension (ACP-CORR-1.0) is planned to specify resource-level aggregated denial signals and group-scoped cooldown as a protocol-level primitive.
Prompt Layer and Jailbreaking. ACP operates at the execution layer, not the prompt layer. Indirect prompt injection [22] and jailbreaking attacks target the model’s inference process before an action request is formed; ACP’s admission boundary is reached only after the agent has produced a concrete action call. These attack classes are therefore outside ACP’s enforcement scope. ACP and prompt-layer defenses are composable: a prompt-layer filter reduces the probability that adversarial instructions generate policy-violating requests; ACP enforces structural admission constraints over the resulting action sequence regardless of how the request was generated. Characterizing the interaction surface between prompt-layer attacks and admission control enforcement is an open research direction [8].
Performance Measurement Scope. All benchmark results reported in Section 13 use InMemoryQuerier, which stores agent state in a process-local hash map protected by a sync.Mutex. This isolates decision-logic latency (739–832 ns p50) from state backend I/O, but does not reflect production latency under persistent backends. A Redis LedgerQuerier introduces one round-trip per evaluation ($\sim$0.1–1 ms on a local network); a PostgreSQL backend introduces 1–5 ms per query depending on index hit rate and connection pool size. Under these conditions the dominant latency is backend I/O, not admission logic, and the 739–832 ns figure becomes a lower bound on end-to-end decision time. The DelayedQuerier experiment (Table 18) characterizes throughput degradation as a function of injected backend latency; benchmarking against a production Redis or PostgreSQL instance is left for future work.
Deterministic Risk Scoring. ACP-RISK-3.0 uses a deterministic, rule-based scoring function. An alternative approach is ML-based anomaly detection (e.g., autoencoders over request sequences or RLHF-trained safety classifiers), which can adapt to distributional shift without explicit rule updates. ACP-RISK is deterministic by design: institutional admission control requires that each decision be independently reproducible and auditable. An ML model cannot produce the fixed, signed admission record required for institutional accountability. BAR-Monitor provides the adaptive detection layer above the deterministic engine: it identifies regime-level shifts in enforcement activity without altering individual admission decisions, preserving auditability while adding adaptive observability.